A cough medicine contains 25% alcohol. How much liquid should a pharmacists add to 120 milliliters of the cough medicine so that it contains only 20% alcohol?
Answer by stanbon(75887)

A cough medicine contains 25% alcohol. How much liquid should a pharmacists add to 120 milliliters of the cough medicine so that it contains only 20% alcohol?
--------------------------
Equation:
alcohol + alcohol = alcohol
0.25*120 + 0*x = 0.20(120+x)
Multiply thru by 100 to get:
25*120 + 0 = 20*120 + 20x
---
20x = 5*120
x = 25 milliliters ( amt. of "liquid" that must be added to the mixture)
================
Cheers,
Stan H.